OpenStudy (anonymous):

Simplify square root of -3. Use complex numbers (i).

OpenStudy (anonymous):

-3=3*(-1) sqrt(3*(-1)=sqrt3 i

OpenStudy (anonymous):

\[\sqrt{-1}=i\] so \[\sqrt{-1*3}=i \sqrt{3}\]

jimthompson5910 (jim_thompson5910):

\[\large \sqrt{-3}=\sqrt{-1}\sqrt{3}=i\sqrt{3}\]

OpenStudy (jwt625):

\[\sqrt{3}i\]